OpenStudy (anonymous):

you have to write 0.62 as a fraction how do you choose your denominator?

OpenStudy (anonymous):

31/50.... use a calculator

myininaya (myininaya):

62/100 which reduces to 31/50

OpenStudy (anonymous):

not solve HOW DID YOU GET YOUR DENOMINATOR

myininaya (myininaya):

the 2 is in the hundredths spot

myininaya (myininaya):

6 is in the tenths spot you could actually write .62=6/10+2/100 then find a common denominator to add the fractions

myininaya (myininaya):

other examples: .53=53/100 .622=622/1000 .6342=6342/10000 .2=2/10 and so on and you can reduce some of these fractions